OpenStudy (anonymous):

24=-6m+1+17 (Show Work)

OpenStudy (anonymous):

24=-6m+18 -18 -18 ----------- 6=6m m=1

OpenStudy (anonymous):

You want to make the 6m onto one side. 24=-6m+1+17 When a number goes onto the other side of the equal sign, then you have to change it's sign. If it was a negative, then it turns into a positive. If it was a positive, then it turns into a negative. 6m=-24+1+17 6m=-6. You want the "m" to be alone, and without that 6. So you divide each side by 6. 6m/6 = -6/6 m=-1